Sheng, Inc. expect to have 21% of return on equity each year. The company's dividend payout ratio is 50%, and it just announced its EPS of $20. Sheng just paid its dividends this year. If Sheng has a beta of 2.1 and the T-bill's return is 10.0%, with investors expect S&P500 to earn a return of 15%.

a. Calculate the intrinsic value of Sheng's common shares. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Intrinsic Value = ?



b-1. What will be the price per share one year from now if Sheng's current market price per share is $109, and people expect that the current market price reflects its intrinsic value? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Price = ?



b-2. What will be your holding-period return if you hold Sheng's common stock for one year? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Expected one-year holding period return = ?
Answer #1

As per CAPM, required return = Risk free rate + beta*(market return – risk free rate)

= 10% + 2.1*(15%-10%)

= 20.5%

Growth rate = ROE*Retention Ratio

= 21%*0.5 = 10.5%

Intrinsic Value = Expected Dividend/(Required Return – Growth Rate)

= 20*50%(1+10.5%)/(20.5%-10.5%)

= $110.5

b-1 Price after one year = Price today(1+Required Return) – Dividend after one year

= 109(1.205) – 11.05

= $120.295

b-2 Holding period return = (120.295+11.05-109)/109

= 20.5%